Calculus I, Fall 2011 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 14

Assigned Dec 15, 2011.

1. Section 7.4: Problems: 27, 46, 55, 57.

2. Section 7.8: Read page 499-500, examples 1-3. Then do problems: 3, 4, 10.

3. Section 8.1: Problems: 3, 7, 19, 23, 29, 33, 35 39, 41, 45, 57, 61, 63, 65, 67, 80 (try as many substitutions as time permits, but at least two). Hint for problems 61-66: Multiply appropriate factor on the numerator and denominator simultaneously. For example, sec θ − tan θ for Problem 63.